It us known that A vector + R vector = B vector and R vector perpendicular A vector .The angle between A vector and B vector is

Grade:12th pass

1 Answers

Yash
2 years ago
Correct option is B)

ย 

2Bโ€‹=A2+B2+2ABcosฮธโ€‹ย ....(i)
โˆดtan90o=A+BcosฮธBsinฮธโ€‹โ‡’A+Bcosฮธ=0
โˆดcosฮธ=โˆ’BAโ€‹
Hence, from (i)ย 4B2โ€‹=A2+B2โˆ’2A2โ‡’A=3โ€‹2Bโ€‹
โ‡’cosฮธ=โˆ’BAโ€‹=โˆ’23โ€‹โ€‹โˆดฮธ=150o